Three times the greater of two consecutive odd integers is five less than four times the smaller. Find the two numbers.
Answer by longjonsilver(2297)

Define:
smallest odd integer = x
larger odd integer = y
we know y - x = 2 eg 19-17=2
and 3y = 4x - 5
so, getting rid of the y (from y=x+2), gives 3(x+2)=4x-5
3x+6=4x-5
6 = x-5
x=11
so y=13
Once you have the answer, just work through the question again with these, to see if it all makes sense.
jon.